XioGonz:

Very simple math q. 6÷2(5+8) = ?

XioGonz:

Simple rule of PEDAS P=( ) E=X^2 M=× D=÷ A=+ S=-

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Simple rule of PEDAS P=( ) E=X^2 M=× D=÷ A=+ S=- \(\color{#0cbb34}{\text{End of Quote}}\) you missed the M

XioGonz:

I made a mistake in the word. Pemdas. ;-;

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Simple rule of PEDAS P=( ) E=X^2 M=× D=÷ A=+ S=- \(\color{#0cbb34}{\text{End of Quote}}\) you missed the M \(\color{#0cbb34}{\text{End of Quote}}\) Yes I'm aware I was trying to delete it and renew it but unfortunately I am unable to

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Simple rule of PEMDAS P=( ) E=X^2 M=× D=÷ A=+ S=- \(\color{#0cbb34}{\text{End of Quote}}\) you missed the M \(\color{#0cbb34}{\text{End of Quote}}\) Yes I'm aware I was trying to delete it and renew it but unfortunately I am unable to \(\color{#0cbb34}{\text{End of Quote}}\) try it hence ---

jhonyy9:

and that s ALL

XioGonz:

I already know how to solve the problem. It's just a controversial question because many people mess up the order and end up getting a different answer.

jhonyy9:

ok but i talk about this missed M from the word PEMDAS

XioGonz:

Yes I know I missed the m. I can't fix it either. I already tried.

Mikaela1509:

the answer 23 it's pretty simple tbh and good day/night

jhonyy9:

but you can re-edit it just click on Quote and edit it and re- post it corrected the wrong posted you can delete it

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mikaela1509 the answer 23 it's pretty simple tbh and good day/night \(\color{#0cbb34}{\text{End of Quote}}\) Can you explain how you got this answer?

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 but you can re-edit it just click on Quote and edit it and re- post it corrected the wrong posted you can delete it \(\color{#0cbb34}{\text{End of Quote}}\) Yeah sure I can quote it, bit as I said earlier, I am unable to delete posts because I'm on my phone. There are certain limitations.

jhonyy9:

ohhh so sorry ok i understand it now

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Simple rule of PEMDAS P=( ) E=X^2 M=× D=÷ A=+ S=-

jhonyy9:

yesssss

jhonyy9:

but 23 not is correct answer on your example

jhonyy9:

@Mikaela1509

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 but 23 not is correct answer on your example \(\color{#0cbb34}{\text{End of Quote}}\) I know. I'm presuming she set it up incorrectly.

Mikaela1509:

no i didn't

XioGonz:

Can you please show your work so that I can walk you through the steps?

Mikaela1509:

I do not need you to walk me through the steps I'm pretty smart when it comes to math and so yeah as i said before good day/night

jhonyy9:

@Mikaela1509 pls understand 23 not is correct answer - we like help you understand it

XioGonz:

I am not trying to make you feel bad and I'm not telling you that you aren't smart. I'm sure you have a reasonable explanation as to why you got your answer. I'm simply letting you know that it's incorrect and I just want to show you the correct format and answer.

Mikaela1509:

ok imma go I need sleep that's probably why I got the answer I did but still makes sense since I haven't had sleep in like a week so yeah

XioGonz:

Ah I see, well I'll solve it below. Good night.

jhonyy9:

good night

Mikaela1509:

oh and btw i don't like being corrected and gn

XioGonz:

Many people don't like being corrected, but that's the way of life. Everywhere you go there will always be somebody correcting you. I assure you it's for the best, correction is key to success.

jhonyy9:

@Mikaela1509 5+8= ?

Mikaela1509:

5+8=13 and 6 divided by 2 = 3

XioGonz:

Yes correct. So now you add 13 and 3

Mikaela1509:

yeah i need sleep and probably my meds too

XioGonz:

Oh. Maybe. So, 13+3= ?

Mikaela1509:

16 ok it's 16 now let me sleep!

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mikaela1509 16 ok it's 16 now let me sleep! \(\color{#0cbb34}{\text{End of Quote}}\) Yes correct! Great job. Good night.

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Yes correct. So now you add 13 and 3 \(\color{#0cbb34}{\text{End of Quote}}\) why you wan adding ? between 2 and parentheses there is multiplication - i dont see there ,,+"

XioGonz:

I forgot to place in the + sign. Let me correct it

jhonyy9:

ohhh !!!!

XioGonz:

Thank you for letting me know jhony! 6÷2+(5+8) 3+13 =16

jhonyy9:

look s good - correct now

XioGonz:

Good job both. Lol.😋

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz Thank you for letting me know jhony! 6÷2+(5+8) 3+13 =16 \(\color{#0cbb34}{\text{End of Quote}}\) np - anytime ...